Brandonsmith:

Erin is a dog groomer she charges $6.00 base fee

AngeI:

Hello, I found the whole question Erin is a dog groomer. She charges a $6.00 base fee plus an extra $0.50 per pound. If x represents the weight of the dog in pounds, which equation models her income for each dog groomed? Income=0.5x+6 income=5x+6 income=6x+0.5 income=6x+5 Since you would be multiplying the $.50 per pound then it would be .5x You would then add that to the 6 dollar base fee to get the total